OpenStudy (anonymous):

Solve the following quadratic equations. (hint: use factorization) HELP ME QUICK! please (2m+3)(4m+3)=0

OpenStudy (radar):

This is already factored. Just set each factor to zero and solve 2m+3=0 2m=-3 m=-3/2 that was one you do the other factor 4m+3

jhonyy9 (jhonyy9):

4m+3=0 4m=-3 m=-3/4

OpenStudy (anonymous):

How about this one? n^2+3n-12=6 Help needed, please?

OpenStudy (precal):

n^2+3n-18=0